OpenStudy (anonymous):

Roger is 5 feet tall and casts a shadow 3.5 feet long. At the same time, the flagpole outside his school casts a shadow 14 feet long. Write and solve a proportion to find the height of the flagpole.

OpenStudy (anonymous):

10 foot

OpenStudy (raden):

5/3.5 = x/14 solve for x (assumed x is the height of the flagpole)

OpenStudy (anonymous):

How

OpenStudy (raden):

look at the left side, 5/3.5 multiply by 4/4, it would becomes 5/3.5 * 4/4 = 20/14, agree ?

OpenStudy (anonymous):

I'm sorry I'm confused I still not sure what the answer is

OpenStudy (raden):

well, look the iluetrasion bellow